### v1

# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

Abstract
Reconstructing human dynamic vision from brain activity is a challenging task with great scientific significance. The difficulty stems from two primary issues: (1) vision-processing mechanisms in the brain are highly intricate and not fully revealed, making it challenging to directly learn a mapping between fMRI and video; (2) the temporal resolution of fMRI is significantly lower than that of natural videos. To overcome these issues, this paper propose a two-stage model named Mind-Animator, which achieves state-of-the-art performance on three public datasets. Specifically, during the fMRI-to-feature stage, we decouple semantic, structural, and motion features from fMRI through fMRI-vision-language tri-modal contrastive learning and sparse causal attention. In the feature-to-video stage, these features are merged to videos by an inflated Stable Diffusion. We substantiate that the reconstructed video dynamics are indeed derived from fMRI, rather than hallucinations of the generative model, through permutation tests. Additionally, the visualization of voxel-wise and ROI-wise importance maps confirms the neurobiological interpretability of our model.
1 Introduction
Researchers in computational neuroscience and the field of artificial intelligence have long sought to decipher and simulate the brain’s visual information processing mechanisms to advance the development of brain-inspired models [1, 2, 3].
In recent years, functional magnetic resonance imaging (fMRI) has emerged as a reliable tool for measuring brain activity due to its high spatial resolution as a non-invasive brain signal recording technique [4]. fMRI-based neural decoders, which map brain signals to visual stimuli, facilitate a deeper understanding of the human visual perception system. Neural decoding can be categorized into classification, identification, and reconstruction, with this study focusing on the most challenging aspect: reconstruction. Prior research has made significant strides in the classification [3, 5, 6, 7, 8] and identification [4, 9, 10, 11] of static stimulus images, and remarkably, some researchers have advanced to the point where they can reconstruct [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] images from brain signals that closely resemble the original stimulus images.
In reality, the majority of visual stimuli we encounter in daily life are continuous and dynamic. As depicted in Figure 1, when a subject views dynamic scenes, the primary visual cortex firstly processes low-level structural information like location, shape, size, and color [23], leading to the preliminary recognition of a black silhouette at the edge of a yellow background. Subsequently, motion information of the object is perceived [24], noting that the silhouette is moving from right to left. Lastly, in the higher visual cortex, the interpretation of category and description gives rise to high-level semantic understanding [25], comprehending the scene as a soldier walking from right to left in a desert.
Figure 1: The human brain’s comprehension of dynamic visual scenes. When receiving dynamic visual information, human brain gradually comprehends low-level structural details such as position, shape and color in the primary visual cortex, discerns motion information, and ultimately constructs high-level semantic information in the higher visual cortex, such as an overall description of the scene.
Due to the inherent nature of fMRI, which relies on the slow blood oxygenation level dependent (BOLD) [26, 27]signal, the sampling frequency is restricted to around 0.5Hz. This frequency is notably lower than the typical 30Hz frame rate of most videos. As a result, a significant discrepancy exists between the temporal resolution of fMRI and the nature video. In fact, each fMRI signal integrates information from approximately 60 video frames. This disparity makes the task of reconstructing video from fMRI signals an exceedingly complex challenge.
To address this challenge, Nishimoto [28] transforms the video reconstruction task into a identification problem, employing the Motion-Energy model [29] and Bayesian inference to reconstruct videos from a predefined video library. Subsequently, Han [30] and Wen [31] et al. map brain responses to the feature spaces of deep neural network (DNN) to reconstruct down-sampled (with the frame rate reduced to 1Hz) video stimuli. Wang [32] et al. develope an f-CVGAN that learns temporal and spatial information in fMRI through separate discriminators [33]. To mitigate the scarcity of video-fMRI data, Kupershmidt [34] et al. utilize self-supervised learning [35] to incorporate a large amount of unpaired video data. These efforts have validated the feasibility of video reconstruction from fMRI, albeit with a lack of explicit semantic information in the results. Chen [36] et al. utilize contrastive learning to map fMRI to the Contrastive Language-Image Pre-Training (CLIP) [37] representation space and fine-tuned inflated Stable Diffusion [38, 39] on a video-text dataset as a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, this work does not consider structure information such as color and position, and it is uncertain whether the motion information in the reconstructed videos originated from the fMRI or the video generation model.
Method Semantic Structure Motion Frame rate Resolution
Nishimoto [28] (Current Biology 2011) × × —— ——
Wen [31] (Cerebral Cortex 2017)
× × —— 64x64
Han [30] (NeuroImage 2019)
× × —— 128x128
Kupershmidt [34] × (✓) 4Hz 112x112
Wang [32] (Cerebral Cortex 2022)
× 4Hz 64x64
Chen [36] (NeurIPS 2023 Oral)
× (✓) 3Hz 256x256
Ours 4Hz 512x512
Table 1: Comparison of modal information used in Mind-Animator and related works. Parentheses indicate the utilization of external video data in the decoding of this feature.
In summary, current video reconstruction models face two challenges:
  • (1)
    As shown in Table 1, they fail to simultaneously capture semantic, structure, and motion information within the reconstructed videos. Moreover, the resolution of the video is low.
  • (2)
    The reliance on external video datasets and video generation models introduces uncertainty regarding the true source of motion information, leading to the possibility that the reconstructed videos may represent a "hallucination" of the video generation model rather than an accurate dynamic decoding from the fMRI data.
To address the aforementioned issues, we introduce Mind-Animator, a video reconstruction model that decouples semantic, structure, and motion information from fMRI, as illustrated in Figure 2. Specifically, we map fMRI to the CLIP representation space and the Vector Quantized-Variational Autoencoder (VQ-VAE) [40] latent space to capture semantic and structure information. We design a Transformer-based [41] motion decoder to extract motion information frame by frame from fMRI through a next-frame-prediction task. Finally, the decoded semantic, structure, and motion information is fed into an inflated Stable Diffusion [38, 39] without any fine-tuning to generate each frame of the video, ensuring that all information is derived solely from the fMRI data.
The contributions are summarized as follows:
  • (1)
    We propose Mind-Animator, which for the first time successfully decouples semantic, structure, and motion information from fMRI to enable video reconstruction. To extract the motion and spatial information from fMRI, we propose temporal and spatial attention modules respectively, which decode subtle but significant motion information.
  • (2)
    We validate through a permutation test that the motion information in our reconstructed videos indeed originates from the fMRI, rather than being a "hallucination" generated by the video generation model.
  • (3)
    We introduce seven evaluation metrics that comprehensively assess the reconstruction results of our model and all previous models across three dimensions—semantic, structure, and spatiotemporal consistency—on three publicly available high-quality video-fMRI datasets. Our model achieves state-of-the-art (SOTA) performance in five of these metrics and secures second place in the remaining two, with a notable 76% improvement in Structural Similarity Index (SSIM) over the previous SOTA. This establishes our work as the first comprehensive and unbiased benchmark for subsequent researchers. The code and data have been anonymously released at: https://github.com/Zuskd/Mind-Animator.
2 Methodology
Figure 2: The overall architecture of Mind-Animator, a two-stage video reconstruction model based on fMRI. As illustrated in subfigures (a), (b), and (c), three decoders are trained during the fMRI-to-feature stage to disentangle semantic, structural, and motion information from fMRI, respectively. Subfigure (d) demonstrates that, in the feature-to-video stage, the decoded information is input into an inflated Text-to-Image (T2I) model for video reconstruction.
2.1 Overview
Figure 2 presents the overall architecture of the proposed Mind-Animator, a video reconstruction model based on fMRI. The model consists of two stages: fMRI-to-feature and feature-to-video.
In the fMRI-to-feature stage, as depicted in Figure 1, we begin by emulating the human visual system’s approach to interpreting dynamic visual stimuli. This process involves the decomposition of video stimuli into high-level semantic feature, low-level structural feature, and motion feature. Then three separate decoders are trained to decode these features from fMRI: (a) for decoding semantic feature, we employ a contrastive learning loss to map fMRI into the visual-linguistic embedding space of CLIP[37], (b) we utilize the frame token extracted by VQ-VAE[40] as the video’s structural feature[42], followed by a simple Multi-Layer Perceptron (MLP) to fit it, and (c) we design a Transformer-based Consistency Motion Generator for decoding motion information. After training with a next-frame-prediction task, this module sequentially generates each subsequent frame token based on the first frame token decoded in section (b).
In the feature-to-video stage, depicted in Figure 2 (d), the decoded features are input into an inflated Text-to-Image (T2I) model, facilitating the reconstruction of the stimulus video without the interference of external training videos.
2.2 Problem Statement
We aim to decode videos from brain activity recorded with fMRI when healthy participants watch a sequence of natural videos. Let $X$ and $Y$ denote the voxel space and pixel space, respectively. Let $X_{i}\in R^{1\times n}$ be the fMRI signal when a video $V_{i,j}\in R^{1\times 3\times 512\times 512}$ is presented to the participant, where $n$ is the number of fMRI voxels, $j$ is the frame ID of video $i$ and $i\in[1,N]$, $j\in[1,8]$, with $N$ the total number of videos. Let $Z(k)$ denote the feature space, $k\in\{semantic,structure,motion\}$. The goal of fMRI-to-feature stage is to train decoders $D(k):X\rightarrow Z(k)$, and the goal of feature-to-video stage is to construct a video generation model $G:Z(semantic)\times Z(structure)\times Z(motion)\rightarrow Y$, without introducing motion information from external video data.
2.3 fMRI-to-feature Stage
Semantic Decoder
Due to the low signal-to-noise ratio of the fMRI signal $X_{i}$ and the substantial dimension discrepancy with the text condition $c_{i}\in R^{1\times 20\times 768}$, directly learning a mapping between them is prone to overfitting. Considering the robust semantic information embedded in the latent space of CLIP[43], and given that CLIP has been shown to outperform various single-modal DNNs in explaining cortical activity[44, 45], we employ bidirectional InfoNCE loss to align the fMRI with the latent space of CLIP (Vit-B/32)$\in R^{512}$, followed by a two-layer MLP to map it to text condition $c_{i}$,
$L_{BiInfoNCE}=-\frac{1}{B}\sum_{i=1}^{B}\Big{(}log\frac{exp(s(\hat{z}_{i},z_{i})/\tau)}{\sum_{j=1}^{B}exp(s(\hat{z}_{i},z_{j})/\tau)}+log\frac{exp(s(\hat{z}_{i},z_{i})/\tau)}{\sum_{k=1}^{B}exp(s(\hat{z}_{i},z_{k})/\tau)}\Big{)}.$ (1)
where $s$ is the cosine similarity, $z_{i}$ and $\hat{z}_{i}$ are the latent representation from two modalities, $B$ is the batch size, and $\tau$ is a learned temperature parameter. Then, given $f\in R^{B\times 512}$, $v\in R^{B\times 512}$, and $t\in R^{B\times 512}$ as the respective representations of fMRI, video, and text embeddings, the fMRI-vision-language trimodal loss is:
$L_{Semantic}=\alpha\cdot L_{BiInfoNCE}(f,t)+(1-\alpha)\cdot L_{BiInfoNCE}(f,v).$ (2)
Subsequently, to map the fMRI embedding $f_{i}$ to the text condition $c_{i}$ for the purpose of conditioning generative image models, a projection loss is utilized,
$L_{Projection}=\frac{1}{B}\sum_{i=1}^{B}\|MLP(f_{i})-c_{i}\|^{2}_{2}.$ (3)
Finally, we combine the Semantic and Projection losses using tuned hyperparameters $\lambda_{1},\lambda_{2}$,
$L_{Combined}=\lambda_{1}\cdot L_{Semantic}+\lambda_{2}\cdot L_{Projection}.$ (4)
Structure Decoder
For a short video clip, it can be assumed that the low-level information (size, shape, and color) contained in each frame remains largely consistent with that of the first frame. Consequently, we utilize the token extracted from the first frame by VQ-VAE as structural information and train the structural decoder using the standard mean squared error (MSE) loss function. Let $\Phi$ denote the encoder of VQVAE, the structure loss is defined as:
$L_{Structure}=\frac{1}{B}\sum_{i=1}^{B}\|D_{Structure}(f_{i})-\Phi(V_{i,1})\|^{2}_{2}.$ (5)
Consistency Motion Generator
Inspired by natural language processing, we treat each video frame token as a word embedding, and develop an L-layer Transformer-based Consistency Motion Generator. For a more detailed introduction, please refer to Appendix B.1.
Figure 3: The detailed architectural diagram of the Consistency Motion Generator.
In the Temporal Module, visible video frame tokens $\Phi(V_{i})\in R^{m\times d_{token}}$ and positional encoding $E_{pos}\in R^{m\times d_{token}}$ are jointly input into a Sparse Causal Self-Attention (SCSA) layer to learn inter-frame temporal information. This attention layer incorporates a specially designed Sparse Causal Mask to ensure sparsity between frames and accelerate training. As illustrated in Figure 3, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts[52]. During inference, we eliminate the random mask. As shown in Eq. 2.3, the model also applies residual connections and layer normalization (LN) to the variable $z_{l}$,
$z_{0}=$ $[\Phi(V_{i,1}),\Phi(V_{i,2}),\ldots,\Phi(V_{i,m})]+E_{pos},$
$z_{l}=$ $LN(SCSA(z_{l-1}))+z_{l-1}.\quad l=1,2,\ldots,L$ (6)
As shown in Eq. 7, in the Spatial Module, the embedding of the visible frames $z_{l}$ serves as the Query, while the fMRI signal $f$, after passing through an embedding layer, serves as the Key and Value in the cross-attention block. Following residual connections and layer normalization, $z_{l}$ is input into the Feed Forward Network (FFN) to predict the subsequent unseen frame tokens $\hat{\Phi(V_{i,j})},j\in[m+1,8]$:
$z_{l}=$ $CrossAttention(Q,K,V),\quad l=1,2,\ldots,L$ (7)
$Q=$ $W_{Q}^{l}\cdot z_{l},\quad K=W_{K}^{l}\cdot Emb(f),\quad V=W_{V}^{l}\cdot Emb(f),$
$z_{l}=$ $FFN(LN(z_{l})+z_{l-1}).\quad l=1,2,\ldots,L$ (8)
Then, the final motion consistency loss is defined as:
$L_{Consistency}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=m+1}^{8}\|\hat{\Phi(V_{i,j})}-\Phi(V_{i,j})\|^{2}_{2}.$ (9)
2.4 Feature-to-video Stage
Inflated Stable Diffusion for Video Reconstruction
Despite the rapid development of video generation models capable of producing vivid videos from text conditions, it is crucial to emphasize that the objective of our project is to disentangle semantic, structural, and motion information from fMRI to fully reconstruct the stimulus video. Utilizing pre-trained video generation models could obscure whether the motion information in the reconstructed video originates from the fMRI or external video data.
To address this issue, we employ the network inflation[46, 47, 39] technique to implement an inflated Stable Diffusion, which is used to reconstruct each frame of the video without introducing additional motion information. For further details, please refer to the Appendix B.2.
3 Experiment
3.1 Datasets
In this study, we utilize three publicly available video-fMRI datasets, which encompass paired stimulus videos and their corresponding fMRI responses. As depicted in Table 2, these datasets collectively comprise brain signals recorded from multiple healthy subjects while they are viewing the videos. The video stimuli are diverse, covering animals, humans, and natural scenery. For detailed information on the datasets and preprocessing steps, please refer to Appendix C.
Dataset Adopted participants TR Train samples Test samples
CC2017[31] 3 2s 4320 1200
HCP[48] 3 1s 2736 304
Algonauts2021[49] 10 1.75s 900 100
Table 2: Characteristics of the video-fMRI datasets used in our experiments
3.2 Evaluation Metrics
To comprehensively and fairly evaluate the performance of our model, we propose the following evaluation metrics.
Semantic-level metrics
Following prior studies[19, 36], we use the N-way top-K accuracy classification test and VIFI-score as the semantics-level metrics. For the classification test, we compare the ground truth (GT) against the predicted video (PV) classifications using a classifier. A trial is successful if the GT class ranks within the top-K probabilities from the PV’s classification among N randomly selected classes. Additionally, we implement two modes: image-based (2-way-I) and video-based (2-way-V). We describe this evaluation method in Algorithm 2. For the VIFI-score, we utilize VIFICLIP[53]—a CLIP model fine-tuned on the video dataset—to extract features from both the GT and the PV, followed by the calculation of cosine similarity.
Pixel-level metrics
We employ the structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and hue-based Pearson correlation coefficient (Hue-pcc) as pixel-level metrics.
Spaciotemporal (ST) -level metric
We adopt the CLIP-pcc, a common metric in the field of video editing, which involves computing CLIP image embeddings on each frame of the predicted videos and reporting the average cosine similarity between all pairs of adjacent video frames.
4 Results
Figure 4: Reconstruction results of CC2017 dataset. Our reconstructed results are highlighted with red boxes, while those of Wen and Nishimoto are delineated by blue and green boxes, respectively.
4.1 Comparative Experimental Results
We compare our model with all previous video reconstruction models [^1] It should be noted that when replicating the results of Nishimoto et al.[28] on the CC2017 dataset, we utilized videos from the training sets of both CC2017 and HCP as the natural movie prior. on the CC2017 dataset. Visual comparisons are presented in Figure 4, while quantitative comparisons are detailed in Table 3. In the computation of quantitative metrics, the results of Wen et al.[31] pertain to the first segment of the test set, whereas the results of other researchers are derived from the whole test set. The findings on HCP and Algonauts2021 datasets are elaborated in Appendices E.2 and E.3, respectively.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.727±$\pm 0.04$ —— —— 0.116±$\pm 0.09$ 8.012±$\pm 2.31$ 0.753±$\pm 0.12$ ——
Wen[31] 0.758±$\pm 0.03$ —— —— 0.114±$\pm 0.15$ 7.646±$\pm 3.48$ 0.647±$\pm 0.11$ ——
Wang[32] 0.713±$\pm 0.04$ 0.773±$\pm 0.03$ 0.596±$\pm 0.07$ 0.118±$\pm 0.08$ 11.432±$\pm 2.42$ 0.589±$\pm 0.18$ 0.402±$\pm 0.41$
Kupershmidt[34] 0.764±$\pm 0.03$ 0.771±$\pm 0.03$ 0.585±$\pm 0.08$ 0.135±$\pm 0.08$ 8.761±$\pm 2.22$ 0.606±$\pm 0.14$ 0.386±$\pm 0.47$
Chen[36] 0.792±$\pm 0.03$ 0.853±$\pm 0.03$ 0.587±$\pm 0.08$ 0.171±$\pm 0.08$ 8.662±$\pm 1.52$ 0.760±$\pm 0.10$ 0.408±$\pm 0.46$
Ours(sub1) 0.809±$\pm 0.03$ 0.837±$\pm 0.02$ 0.602±$\pm 0.07$ 0.301±$\pm 0.09$ 9.134±$\pm 1.48$ 0.768±$\pm 0.12$ $\pm 0.42$
Ours(sub2) 0.804±$\pm 0.29$ 0.832±$\pm 0.03$ 0.604±$\pm 0.08$ 0.287±$\pm 0.11$ 9.049±$\pm 1.45$ 0.795±$\pm 0.12$ $\pm 0.42$
Ours(sub3) 0.792±$\pm 0.03$ 0.833±$\pm 0.03$ 0.600±$\pm 0.08$ 0.349±$\pm 0.11$ 9.306±$\pm 1.54$ 0.791±$\pm 0.12$ 0.415±$\pm 0.39$
Table 3: Quantitative comparison of reconstruction results. All metrics indicate superior performance with larger values, with the best results highlighted in bold and the second-best results underlined.
Table 3 indicates that our model achieves SOTA performance in 5 out of 7 metrics, securing the second place in the remaining two. Specifically, our model outperforms the previous SOTA model by 76% in terms of SSIM, which underscores the benefits of incorporating structural information. Specifically, as depicted in Figure 4, our reconstruction results contain richer semantic information compared to earlier models, such as a girl and a yellow dog being held in someone’s arms. In contrast to Mind-video by Chen et al.[36], our results are more consistent with the ground truth in terms of fine-grained structural and motion information. For instance, the reconstructed girl’s clothing color, the dog’s fur color, and the positioning of the forest along the coastline are closer to the stimulus videos. Regarding motion information, our results depict the dog being petted and a noticeable camera movement in the coral reef scene. Additional reconstruction results on other subjects, as well as instances of reconstruction failure, are presented in Appendix E.1.
4.2 Ablation Study
Semantic-$\uparrow$ Pixel-$\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
w/o Semantic 0.679±$\pm 0.04$ 0.766±$\pm 0.04$ 0.523±$\pm 0.07$ $\pm 0.09$ 8.005±$\pm 1.57$ 0.737±$\pm 0.11$ 0.123±$\pm 0.31$
w/o Structure 0.789±$\pm 0.03$ 0.814±$\pm 0.03$ 0.555±$\pm 0.08$ 0.184±$\pm 0.08$ 8.712±$\pm 1.37$ 0.791±$\pm 0.11$ 0.260±$\pm 0.41$
w/o Motion 0.674±$\pm 0.04$ 0.789±$\pm 0.03$ 0.585±$\pm 0.08$ $\pm 0.13$ 8.611±$\pm 2.43$ 0.715±$\pm 0.14$ 0.376±$\pm 0.42$
Full Model 0.809±$\pm 0.03$ 0.837±$\pm 0.02$ 0.602±$\pm 0.07$ 0.301±$\pm 0.10$ 9.134±$\pm 1.51$ 0.768±$\pm 0.11$ 0.425±$\pm 0.41$
Table 4: Ablation study on our proposed decoders. 100 repetitions are conducted on the metrics 2-way-I and 2-way-V, while 5 trials are performed on other metrics, with the results being averaged across all samples in test set and trials. Colors reflect statistical significance (paired t-test) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
In this subsection, we conduct a detailed ablation study to assess the effectiveness of the three decoders we proposed and to evaluate the impact of various hyperparameters on video reconstruction (See Appendix E.4).
First, we present the results obtained using the full model. Then, on the basis of the full model, we separately eliminate the semantic decoder (w/o Semantic) and the structure decoder (w/o Structure) by replacing their outputs with random noise. For the consistency motion generator, we replaced it with 8 simple MLPs to model each frame individually (w/o Motion). Table 4 demonstrates that the removal of any decoder results in a significant decline in the model’s performance across nearly all metrics, which shows the efficacy of our proposed decoders.
5 Interpretability Analysis
5.1 Have we truly decoded motion information from fMRI?
This work focuses on the video reconstruction from fMRI, aiming for motion consistency between the reconstructed and stimulus videos. We specifically design a Consistency Motion Generator (CMG) to decode motion information from fMRI. Following the work of Wang et al. [32], we perform a permutation test on 3 subjects from the CC2017 dataset to ascertain whether this module decodes the correct motion information from fMRI. Specifically, for each 8-frame reconstructed video clip from each subject, we randomly shuffle the frame order 100 times and compute pixel-level and spaciotemporal-level metrics between the actual and shuffled frames. Subsequently, we estimate the P-value by the following formula: $P=\sum_{i=1}^{100}\delta_{i}/100$, where $\delta_{i}=1$ if the $i$th permutation outperforms the reconstruction result in the original order based on the metrics; otherwise, $\delta_{i}=0$. A lower P-value signifies a closer alignment between the sequential order of the reconstructed video and the ground truth. We repeat the permutation test 5 times under conditions with and without the CMG, as illustrated in Figure 5. It can be observed that the P-value significantly increased across nearly all metrics for all subjects when the CMG is removed, suggesting that we truly decodes motion information from fMRI.
Figure 5: The result of permutation test on the CC2017 dataset. The experiment is repeated 5 times on 3 subjects, with the mean and std presented in subplots (a), (b), and (c), respectively. Paired t-tests are performed, with significance denoted as $p<0.001(***)$, $p<0.01(**)$, $p<0.05(*)$, and $p>0.05(NS)$ for non-significant results.
5.2 Which brain regions are responsible for decoding different features, respectively?
To investigate voxels in which brain regions are responsible for decoding different features (semantic, structure, motion) during the fMRI-to-feature stage, we compute the voxel-wise importance maps in the visual cortex. Specifically, for a trained decoder, we multiply the weight matrix of the linear layers, then average the result across the feature dimension, and normalize it to estimate the importance weight for each voxel. A higher weight indicates that the voxel plays a more significant role in feature decoding. We project the importance maps of subject 1’s voxels from the CC2017 dataset onto the visual cortex, as depicted in Figure 6. To obtain ROI-wise importance maps, we calculate the average of the importance weights of voxels contained within each Region of Interest (ROI), with the results presented in Figure 7. The results from other subjects are presented in Appendix E.5.
Figure 6: Voxel-wise importance maps projected onto the visual cortex of subject 1. The lighter the color, the greater the weight of the voxel in the interpretation of feature.
Figure 6 (a) indicates that high-level visual cortex (HVC, such as MT, MST and TPOJ) contribute more significantly to the decoding of semantic feature, with a calculated weight of 2.588, accounting for 60.5% of the total, as shown in Figure 7 (a). In contrast, low-level visual cortex (LVC, such as V1, V2, V3) have a weight of 1.685, representing 39.5%. Although it is not immediately apparent from Figure 6 (b) which ROI contributes most to the decoding of structural feature, Figure 7 (b) reveals that V1 and V2 have the greatest weight, with HVC having a weight of 1.279 (36.03%), and LVC weighing 2.271 (63.97%). Considering the aforementioned findings, our results are plausible from the perspective of cognitive neuroscience. It is generally believed that LVC is predominantly responsible for processing low-level information of visual stimuli [4, 54, 55], such as orientation and contour. Meanwhile, V4 is involved in color processing [56, 57]. In contrast, HVC is responsible for processing high-level semantic information of objects[58], including category.
Figure 7: ROI-wise importance maps in the visual cortex of subject 1.
Figure 6 (c) indicates that both LVC and HVC contribute to the decoding of motion information, with significant weight attributed to V1 and MT. As derived from Figure 7 (c), the weight distribution between LVC and HVC is comparable, accounting for 42.4% and 57.6%, respectively. This observation is consistent with previous work[59], which validates the function of MT in visual motion processing. Furthermore, our findings affirm that the spatial and temporal modules designed in CMG effectively capture spatiotemporal information from across both LVC and HVC.
6 Conclusion
In this paper, we introduce a video reconstruction model (Mind-Animator) that decouples semantic, structural, and motion information from fMRI, achieving state-of-the-art performance across 3 public datasets. We mitigate the interference of external video data on motion information decoding through a rational experimental design. The results of the permutation test demonstrate that the motion information we decoded indeed originates from fMRI, rather than being a "hallucination" from generative model. Additionally, the visualization of voxel-wise and ROI-wise importance maps substantiate the neurobiological interpretability of our model.
Acknowledgments and Disclosure of Funding
We would like to express our gratitude to Prof.Jack L. Gallant and Prof.Shinji Nishimoto for their pioneering exploration in the field of video reconstruction and for their high-quality code. We are grateful to Prof.Juan Helen Zhou and Dr.Zijiao Chen for their patient answers to our questions and for making all the results of the Mind-video test set public. We also extend our thanks to Prof.Michal Irani, Dr.Ganit Kupershmidt, and Dr.Roman Beliy for providing us with all the reconstruction results of their models on the test set.
We would like to express our appreciation to Prof.Zhongming Liu and Dr.Haiguang Wen for their open-sourced high-quality video-fMRI dataset and the preprocessing procedures. Our gratitude also goes to the Human Connectome Project (HCP) for providing a large-scale fMRI dataset and cortical visualization tools. We are thankful to the Algonauts2021 competition for providing a set of pre-processed video-fMRI data from multiple subjects.
We are thankful to the Stable Diffusion team for their high-performance text-to-image model, and we also appreciate the Tune-a-video team for their open-source video editing framework, which allows us to reconstruct videos without introducing additional motion information.
Appendix A Related Work
A.1 Reconstructing Human Vision from Brain Activities
Reconstructing Images from Brain Activities
Building on Haxby’s[60] seminal work, the field of neural decoding has seen a proliferation of tasks with significant implications for guiding research. These tasks can be broadly classified into three categories: stimulus classification, identification, and reconstruction, with the latter being the most challenging and the focus of our study.
Traditional image reconstruction techniques rely on linear regression models to correlate fMRI with manually defined image features[61, 12, 8], yielding blurry results and a heavy reliance on manual feature selection. However, the advent of deep learning has revolutionized this domain. Deep neural networks (DNNs) have become increasingly prevalent for their ability to address the scarcity of stimulus-fMRI pairs through semi-supervised learning[62], as demonstrated by Beliy et al.[22] and Gaziv et al.[63]. Yet, these models often fail to capture discernible semantic information. Chen et al.[19] employed a pre-training and fine-tuning approach on fMRI data, leveraging methods akin to Masked Autoencoder(MAE)[64] and Latent Diffusion Models(LDM)[38] to improve reconstruction quality. Ozcelik et al.[21] and Gu et al.[65] utilized self-supervised models for feature extraction, followed by iterative optimization to refine the reconstruction process. The integration of semantic information from text, facilitated by Contrastive Language-Image Pre-Training (CLIP)[37] , has been instrumental in reconstructing complex natural images. Lin et al.[16] and Takagi et al.[20] demonstrated the potential of aligning fMRI with CLIP representations and mapping fMRI to text and image features for high-fidelity reconstruction.
While rapid advancements have been made in stimulus reconstruction, with some researchers achieving reconstructions from brain signals that closely approximate the original stimuli, the majority of prior work has focused on static image reconstruction. This study, however, shifts the focus to the more challenging task of video reconstruction.
Reconstructing Videos from Brain Activities
Compared to image reconstruction, the challenge in video reconstruction lies in the significant discrepancy between the temporal resolution of fMRI (0.5Hz) and the frame rate of the stimulus video (30Hz), which presents a substantial challenge in modeling the mapping between fMRI signals and video content. To overcome the challenge, Nishimoto [28] transformed the video reconstruction task into a identification problem, employing the Motion-Energy model [29] and Bayesian inference to reconstruct videos from a predefined video library. Subsequently, Han [30] and Wen et al.[31] mapped brain responses to the feature spaces of DNN to reconstruct down-sampled (with the frame rate reduced to 1Hz) video stimuli. The aforementioned studies have preliminarily validated the feasibility of reconstructing video frames from fMRI. Wang et al.[32] developed an f-CVGAN that learns temporal and spatial information in fMRI through separate discriminators [33]. To mitigate the scarcity of fMRI-video data, Kupershmidt et al.[34] utilized self-supervised learning [35] to incorporate a large amount of unpaired video data. These efforts have validated the feasibility of video reconstruction from fMRI, albeit with a lack of explicit semantic information in the results. Chen et al.[36] utilized contrastive learning to map fMRI to the CLIP representation space and fine-tuned inflated Stable Diffusion [38, 39] on a video-text dataset as a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, Chen did not consider structure information such as color and position, and it was uncertain whether the motion information in the reconstructed videos originated from the fMRI or the video generation model.
A.2 Diffusion Models
Diffusion models[66, 67] , a class of probabilistic generative models, have increasingly rivaled or surpassed the performance of Generative Adversarial Networks (GAN)[33] in specific tasks within the field of computer vision. Diffusion models encompass a forward diffusion process and a reverse denoising process, each exhibiting Markovian behavior. The forward process incrementally introduces Gaussian noise into the original image, culminating in a transition to standard Gaussian noise. The forward diffusion process can be represented as $q(x_{t}|x_{t-1})=N(x_{t};\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})I)$, where $t$ denotes the time step of each noise addition. The reverse denoising process employs the U-Net[68] architecture to accurately model the noise distribution at each timestep $t$. The image synthesis is achieved through a sequential denoising and sampling procedure, initiated from standard Gaussian noise.
In the context of image generation tasks, the conventional diffusion model executes two Markov processes in a large pixel space, resulting in substantial computational resource utilization. To address this issue, Latent Diffusion Models (LDM)[38] employs a VQ-VAE[40] encoder to transform the pixel space into a low-dimensional latent space. Subsequently, the diffusion model’s training and generation are performed in the latent space, with the final generated image obtained by utilizing the VQ-VAE decoder. This approach significantly reduces computational resource requirements and inference time while preserving the quality of generated images.
A.3 Diffusion Models for Video Generation
After achieving significant progress in text-to-image (T2I) generation tasks, diffusion models have piqued the interest of researchers in exploring their potential for text-to-video (T2V) generation. The pioneering work by Ho et al.[69] , introducing the 3D diffusion U-Net, marked significant progress in applying Diffusion Models to video generation. This was followed by further advancements by Ho et al.[70] , who utilized a cascaded sampling framework and super-resolution method to generate high-resolution videos. Subsequent contributions have expanded upon this work, notably with the incorporation of a temporal attention mechanism over frames by Singer et al.[71] in Make-A-Video. Zhou et al.[72] with MagicVideo, and He et al.[73] with LVDM, have integrated this mechanism into latent Diffusion Models, significantly enhancing video generation capabilities. However, due to the scarcity of paired text-video datasets and the high memory requirements for training 3D U-Nets, alternative approaches are being explored. These involve refining pre-trained T2I models to directly undertake T2V tasks. Khachatryan et al.[47] introduced two enhancements to enable zero-shot adaptation of T2I models to T2V tasks: (1) the implementation of cross-frame attention, ensuring that the generation of each current frame in a video considers information from preceding frames; and (2) the consideration of inter-frame correlations during noise sampling, rather than random sampling for each frame independently. Wu et al.[39] also employed cross-frame attention and achieved one-shot video editing by fine-tuning partial model parameters on individual videos.
In this work, tasked with video reconstruction from fMRI, we eschewed the use of pre-trained T2V models to mitigate the interference of external video data with the decoding of motion information from fMRI. Inspired by the cross-frame attention mechanism, we adapted a T2I model through network inflation techniques, enabling it to generate multi-frame videos. Consequently, the generative model employed in our study has never been exposed to video data, ensuring that the motion information in the reconstructed videos is solely derived from the fMRI decoding process.
Appendix B Model Architecture
B.1 Consistency Motion Generator
Figure 3 illustrates the Consistency Motion Generator, which is primarily composed of two modules: the Temporal Module and the Spatial Module.
The Temporal Module is tasked with learning the temporal dynamics from the visible frames. Given the severe information redundancy between video frame tokens, we specifically design a Sparse Causal mask. As shown in Figure 3 on the right, during training, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts[52] and accelerating training. During inference, we eliminate the random mask to allow full utilization of information from all preceding frames for predicting future frames.
Since a single fMRI frame captures information from several video frames, we design a cross attention mechanism within the Spatial Module to extract the necessary temporal and spatial information for predicting the next frame token from the fMRI data.
B.2 Text-to-Image Network Inflation
To leverage pre-trained weights from large-scale image datasets, such as ImageNet, for the pre-training of massive video understanding models, Carreira et al.[46] pioneered the expansion of filters and pooling kernels in 2D ConvNets into the third dimension to create 3D filters and pooling kernels. This process transforms N×N filters used for images into N×N×N 3D filters, providing a beneficial starting point for 3D video understanding models by utilizing spatial features learned from large-scale image datasets.
In the field of generative model, several attempts have been made to extend generative image models to video models. A key technique employed in this work involves augmenting the Query, Key, and Value of the attention module, as illustrated below:
$Q=W^{Q}\cdot z_{v_{i}},\quad K=W^{K}\cdot[z_{v_{0}},z_{v_{i-1}}],\quad V=W^{V}\cdot[z_{v_{0}},z_{v_{i-1}}],$ (10)
where $z_{v_{i}}$ denotes the latent of the $i$-th frame during the generation process.
Appendix C Data Preprocessing
For the stimulus videos of the three datasets described below, we segmented them into 2-second clips, down-sampled the frame rate to 4Hz (i.e., evenly extracting 8 frames), then centrally cropped each frame, and resized each to a shape of 512x512. Following the approach of Chen et al.[36], we employed BLIP2[74] to obtain textual descriptions for each video clip, with lengths not exceeding 20 words.
C.1 Video Captioning with BLIP2
PyTorch code for the video captioning process is depicted in Algorithm 1.
Algorithm 1 PyTorch code for the video captioning process
import torch
import clip
import random
import numpy as np
from PIL import Image
from lavis.models import load_model_and_preprocess
device = torch.device(’cuda:5’)
clip_model, clip_preprocess = clip.load("ViT-B/32", device=device)
model,vis_processors,_=load_model_and_preprocess(name="blip2_t5",
model_type="pretrain_flant5xxl", is_eval=True, device=device)
prompt = "Question:Whatdoesthisimagedescribe?Answer:"
def Video_Captioning(Train_video_path_root):
Train_Captions = []
for i in tqdm(range(18)):
for j in tqdm(range(240)):
[…]
frame1 = Image.open(frames_root + ’0000000.jpg’).convert("RGB")
frame1 = vis_processors["eval"](frame1).unsqueeze(0).to(device)
frame2 = Image.open(frames_root + ’0000056.jpg’).convert("RGB")
frame2 = vis_processors["eval"](frame2).unsqueeze(0).to(device)
frame_mid = Image.open(frames_root + ’0000024.jpg’).convert("RGB")
clip_image = clip_preprocess(frame_mid).unsqueeze(0).to(device)
caption1 = model.generate({"image": frame1, "prompt": prompt})
caption2 = model.generate({"image": frame2, "prompt": prompt})
text1 = clip.tokenize(caption1).to(device)
text2 = clip.tokenize(caption2).to(device)
with torch.no_grad():
image_features3 = clip_model.encode_image(clip_image)
text_features1 = clip_model.encode_text(text1)
text_features2 = clip_model.encode_text(text2)
cos_sim1 = torch.cosine_similarity(image_features3, text_features1)
cos_sim2 = torch.cosine_similarity(image_features3, text_features2)
if abs(cos_sim1 - cos_sim2) <= 0.05:
number = random.random()
if number >= 0.5:
caption = caption1[0]
else:
caption = caption2[0]
else:
caption = caption1[0] + ’,andthen + caption2[0]
Train_Captions.append(caption)
Train_Captions = np.array(Train_Captions)
return Train_Captions
C.2 CC2017
CC2017[31] dataset was first used in the work of Wen et al.[31] This dataset include fMRI data from 3 subjects who view a variety of movie clips (23° × 23°) with a central fixation cross (0.8° × 0.8°). Clips are divided into 18 training movies and 5 testing movies, each eight minutes long, presented 2 and 10 times to each subject, respectively. MRI (T1 and T2-weighted) and fMRI data (2-second temporal resolution) are acquired using a 3-T system. The fMRI volumes are processed for artifact removal, motion correction (6 DOF), registered to MNI space, and projected onto cortical surfaces coregistered to a template.
To extract voxels in the activated visual areas, we calculate the correlation of the time series for each voxel’s activation across 2 trials within the training set. Subsequently, we apply Fisher’s z-transform to the computed correlations, average the results across 18 sessions, and identify the most significant 4500 voxels using a paired t-test to form a mask. This mask is computed separately for each of the 3 subjects on their respective training sets and then applied to both training and test set data, with the selected voxels averaged across trials. Following the work of Nishimoto et al.[28] and Han et al. [30] , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay.
C.3 HCP
This dataset is part of the Human Connectome Project (HCP)[48], encompassing BOLD (blood-oxygen-level dependent) responses from 158 subjects. For the subsequent experiments, three subjects (100610, 102816, and 104416) are randomly selected from this dataset. Data acquisition is performed using a 7T MRI scanner with a spatial resolution of 1.6 millimeters and a repetition time (TR) of 1 second. The utilized BOLD signals undergo standard HCP preprocessing procedures, which include correction for head motion and distortion, high-pass filtering, and removal of temporal artifacts via independent component analysis (ICA). The preprocessed BOLD responses are then registered to the MNI standard space.
Due to the difficulty in directly acquiring fMRI data from multiple trials, we directly utilize the parcellation of the human cerebral cortex proposed by Glasser et al.[75] to extract voxels within the activated visual cortex. The resulting ROIs we extract include: V1, V2, V3, hV4, PPA, FFA, LO, PHC, MT, MST, and TPOJ, totaling 5820 voxels. Following the work of Nishimoto et al.[28] and Han et al.[30] , we utilize BOLD signals with a 4-second lag to represent the movie stimulus responses, thereby accounting for the hemodynamic response delay. Given that no prior work has conducted video reconstruction experiments on these three subjects from the HCP dataset, except for Wang et al.[32], we randomly shuffle all video segments and allocate 90% for the training set, with the remaining 10% reserved for the test set.
C.4 Algonauts2021
This dataset is publicly released for the 2021 Algonauts Challenge[49]. During data acquisition, 10 participants passively view 1100 silent videos of everyday events, each approximately 3 seconds in duration, presented three times. The participants’ fMRI is recorded using a 3T Trio Siemens scanner with a spatial resolution of 2.5 millimeters and a repetition time (TR) of 1.75 seconds. The fMRI preprocessing involves steps such as slice-timing correction, realignment, coregistration, and normalization to the MNI space. Additionally, the fMRI data are interpolated to a TR of 2 seconds.
The dataset has been officially preprocessed, allowing us to extract brain responses from nine regions of interest (ROIs) within the visual cortex, including four primary and intermediate visual cortical areas V1, V2, V3, and V4, as well as five higher visual cortical areas: the Extrastriate Body Area (EBA), Fusiform Face Area (FFA), Superior Temporal Sulcus (STS), Lateral Occipital Cortex (LOC), and Parahippocampal Place Area (PPA). These areas selectively respond to body, face, biological motion and facial information, objects, and scene information, respectively. In the experiment, the average neural response across three stimulus repetitions is taken for brain activity. As the test set data are not yet public, we utilize the first 900 sets of data for training and the 900-1000 sets for testing.
C.5 Data Acquisition
The open-source datasets used in this paper can be accessed via the following links:
Appendix D Implementation Details
D.1 Hyperparameter Settings
For all three datasets employed in the experiments, during the training of the Semantic Decoder, we set $\alpha$ to 0.5, $\lambda_{1}$ to 0.01, and $\lambda_{2}$ to 0.5. The batch size is set to 64, and the learning rate is set to 2e-4, with training conducted 100 epochs. Given the critical role of data volume and augmentation methods in the training of contrastive learning models, and the scarcity of video-fMRI paired data, we implement specific data augmentation techniques to prevent overfitting. For fMRI data, we randomly select 20% of the voxels during each iteration, zero out 50% of their values. For image data, we randomly crop one frame from eight video frames to a size of 400x400 pixels, and then resize it to 224x224. For text data, due to the presence of similar video clips in the training set (derived from a complete video segment), BLIP2 often provide identical textual descriptions, which leads to overfitting. To mitigate this, we apply more aggressive augmentation techniques. For an input sentence, we perform Synonym Replacement with a 50% probability and Random Insertion, Random Swap, and Random Deletion with a 20% probability each.
During the training of the Structural Decoder, we set the batch size to 64 and the learning rate to 1e-6. To stabilize the learning process, we conduct the training process for 100 epochs with a 50-step warmup.
When training the Consistency Motion Generator, we set the mask ratio of the Sparse Causal mask to 0.6 during the training phase, with a batch size of 64 and a learning rate of 4e-5. Similarly, for stability, we implement a 50-step warmup followed by 300 epochs of training.
Taking three subjects from the CC2017 dataset as an example, we utilize the first 4000 data points as the training set and the subsequent 320 data points as the validation set. Following this, we retrain the model on the entire training set, which comprises 4320 data points. The loss curve is depicted in Figure 8 .
Figure 8: The training and validation loss curves for subject 1 in the CC2017 dataset.
According to Figure 8, in the inference phase, we utilize model parameters saved at the 30th, 75th, and 70th epochs for the Semantic Decoder, Structural Decoder, and Consistency Motion Generator, respectively. For the generative model, we employ inflated Stable Diffusion V1-5. Given that Stable Diffusion operates on Gaussian-distributed latent space inputs (i.e., $Z_{T}$) for Text-to-Image task, and the distribution of decoded structural information does not align with this, we apply 250 steps of Gaussian smoothing to it, which includes 50 steps of ddim inversion. All experiments are conducted on a single A100 80G GPU, with the training phase taking 8 hours and the inference phase taking 12 hours for each dataset.
D.2 Evaluation Metric Implementation
Semantic-level
This algorithm performs the N-trial n-way top-1 classification test. We describe our evaluation method in Algorithm 2. For the image-based scenario, we utilize a Vision Transformer (ViT)[50] pre-trained on ImageNet as the classifier. For video-based scenario, a pre-trained VideoMAE[52] is employed as the classifier.
Algorithm 2 N-trial n-way top-1 classification test
1:  Input pre-trained classifiers $\mathcal{C}_{image}(\cdot)$, $\mathcal{C}_{video}(\cdot)$, video pair (Generated Video $x$, Corresponding GT Video $\hat{x})$, mode(video-based or image-based)
2:  Output success rate $r\in[0,1]$
3:  if mode=’video-based’ then
4:     for $N$ trials do
5:        $\hat{y}\leftarrow\mathcal{C}_{video}(\hat{x})$ get the ground-truth class
6:        $\{p_{0},...,p_{399}\}\leftarrow\mathcal{C}_{video}(x)$ get the output probabilities
7:        $\{p_{\hat{y}},p_{y_{1}},...,p_{y_{n-1}}\}\leftarrow$ pick $n$-1 random classes
8:        success if $\mathop{\arg\max}\limits_{y}\{p_{\hat{y}},p_{y_{1}},...,p_{y_{n-1}}\}=\hat{y}$
9:     end for
10:     $r=$ number of success / $N$
11:  else
12:     for 8 frames do
13:        for $N$ trials do
14:           $\hat{y}_{i}\leftarrow\mathcal{C}_{image}(\hat{x}_{i})$ get the ground-truth class
15:           $\{p_{0},...,p_{999}\}\leftarrow\mathcal{C}_{image}(x_{i})$ get the output probabilities
16:           $\{p_{\hat{y}_{i}},p_{y_{i,1}},...,p_{y_{i,n-1}}\}\leftarrow$ pick $n$-1 random classes
17:           success if $\mathop{\arg\max}\limits_{y_{i}}\{p_{\hat{y}_{i}},p_{y_{i,1}},...,p_{y_{i,n-1}}\}=\hat{y}_{i}$
18:        end for
19:        $r_{i}=$ number of success / $N$
20:     end for
21:     $r=\sum_{i=1}^{8}r_{i}/8$
22:  end if
Spaciotemporal-level
Considering the input to the video reconstruction task contains substantial noise, there are instances where every pixel of each reconstructed video frame is either zero or noise, which would artificially inflate the CLIP score if used directly. Therefore, we calculate the CLIP score only when the VIFI-CLIP value exceeds 0.6; otherwise, we assign a score of 0.
Appendix E Additional Experimental Results
E.1 Further Results on the CC2017 Dataset
As depicted in Table 5 and 6, we further demonstrate the experimental results on subjects 2 and 3 from the CC2017 dataset. Specifically, for the metrics 2-way-I and 2-way-V, we present the mean and standard deviation across all samples over 100 trials, for the remaining metrics, we offer the mean and standard deviation of all results within the test set to reflect the overall performance in reconstruction.
E.1.1 CC2017 sub2
As shown in Table 5, our model on subject 2 achieves SOTA performance in five out of seven metrics and secured second place in one.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.787±$\pm 0.04$ —— —— 0.112±$\pm 0.07$ 8.592±$\pm 2.11$ 0.713±$\pm 0.11$ ——
Wen[31] 0.783±$\pm 0.03$ —— —— $\pm 0.14$ 8.415±$\pm 3.46$ 0.626±$\pm 0.12$ ——
Wang[32] 0.727±$\pm 0.04$ 0.779±$\pm 0.03$ 0.596±$\pm 0.07$ 0.107±$\pm 0.06$ 10.940±$\pm 1.83$ 0.589±$\pm 0.19$ 0.404±$\pm 0.38$
Kupershmidt[34] 0.776±$\pm 0.03$ 0.766±$\pm 0.03$ 0.591±$\pm 0.08$ 0.157±$\pm 0.09$ 11.914±$\pm 1.89$ 0.601±$\pm 0.18$ 0.382±$\pm 0.41$
Chen[36] 0.789±$\pm 0.03$ 0.842±$\pm 0.03$ 0.595±$\pm 0.08$ 0.172±$\pm 0.09$ 8.929±$\pm 1.62$ 0.773±$\pm 1.13$ 0.409±$\pm 0.46$
Ours(sub2) 0.804±$\pm 0.03$ 0.832±$\pm 0.03$ 0.604±$\pm 0.08$ 0.287±$\pm 0.11$ 9.049±$\pm 1.45$ 0.795±$\pm 0.12$ $\pm 0.42$
Table 5: Quantitative comparison of reconstruction results on subjects 2.
E.1.2 CC2017 sub3
As shown in Table 6, our model on subject 3 achieves SOTA performance in four out of seven metrics and secured second place in two. It should be noted that the primary objective of this work is to attempt the disentanglement of semantic, structural, and motion information from fMRI data, rather than achieving SOTA performance. The superior results obtained in the aforementioned metrics can more compellingly demonstrate the effectiveness of our proposed model.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.712±$\pm 0.04$ —— —— 0.128±$\pm 0.07$ 8.546±$\pm 1.91$ 0.746±$\pm 0.12$ ——
Wen[31] —— —— —— —— —— —— ——
Wang[32] 0.722±$\pm 0.04$ 0.778±$\pm 0.03$ 0.584±$\pm 0.07$ 0.098±$\pm 0.06$ 10.758±$\pm 2.00$ 0.572±$\pm 0.49$ 0.392±$\pm 0.41$
Kupershmidt[34] 0.767±$\pm 0.03$ 0.766±$\pm 0.03$ 0.597±$\pm 0.08$ 0.128±$\pm 0.08$ 11.237±$\pm 1.86$ 0.641±$\pm 0.14$ $\pm 0.37$
Chen[36] 0.811±$\pm 0.03$ 0.848±$\pm 0.03$ 0.597±$\pm 0.08$ 0.187±$\pm 0.10$ 9.013±$\pm 1.66$ 0.771±$\pm 0.13$ 0.410±$\pm 0.46$
Ours(sub3) 0.792±$\pm 0.03$ 0.833±$\pm 0.02$ 0.600±$\pm 0.07$ 0.349±$\pm 0.11$ 9.306±$\pm 1.54$ 0.791±$\pm 0.12$ 0.415±$\pm 0.39$
Table 6: Quantitative comparison of reconstruction results on subjects 3.
E.1.3 More Reconstruction Results on Multiple Subjects
Due to the anatomical and functional connectivity differences among subjects, even when presented with the same stimulus video, different brain signals are elicited. Consequently, we train the Mind-Animator on three subjects from the CC2017 dataset separately, with the reconstruction results shown in Figure 9. It can be observed that, despite training and applying the model directly to the three subjects without additional modifications, the reconstruction outcomes are largely consistent, which substantiates the effectiveness of our model.
However, as it is challenging to obtain a large volume of brain signals from the same subject, the direct extension of a model trained on one subject to others, or the direct training using data from multiple subjects, represents a future research direction that requires further investigation.
Figure 9: The reconstruction results on three subjects from the CC2017 dataset.
Figure 10: More reconstruction results on the CC2017 dataset.
Figure 11: Reconstruction failure cases.
In Figure 10, we present additional reconstruction results from the CC2017 dataset, demonstrating that our model does not overfit despite the limited data volume. It is capable of decoding a rich array of video clips, such as two people conversing, an airplane in flight, and a dog turning its head, among others. These video clips encompass a wide range of natural scenarios found in everyday life, including human activities, animal behaviors, and natural landscapes.
Additionally, to provide a comprehensive and objective assessment of our model, we also include some reconstruction failure cases in Figure 11. The primary reasons for these failures are twofold. Firstly, the inherent data acquisition paradigm leads to abrupt transitions in content at the junction of video clips, which are evenly segmented from complete videos watched by the subjects during data collection. As illustrated in the first row of Figure 11, the model often struggles to recognize such abrupt changes, resulting in the reconstruction of only the scene prior to the transition. Secondly, errors in feature decoding occur, as shown in the second row of Figure 11. Although the model successfully decodes an ocean scene, structural decoding errors led to the appearance of numerous fish in the reconstructed frames.
E.2 Further Results on the HCP Dataset
To validate the broad effectiveness of our proposed model across different datasets, we also conduct experiments on the HCP dataset. The selection of three subjects (100610, 102816, and 104416) aligns with that of Wang et al.[32], and the parameter settings for Mind-Animator during both training and testing are identical to those used on the CC2017 dataset. We replicate Nishimoto et al.[28] ’s model on HCP, utilizing video clips from the training sets of both CC2017 and HCP as video priors. Wang only reports SSIM and PSNR metrics on the HCP dataset. Visual comparisons are presented in Figure 12, while quantitative comparisons are detailed in Table 7. The results indicate that our model still performs well across multiple subjects in the HCP dataset.
Figure 12: The reconstruction results on three subjects from the HCP dataset.
Sub ID Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Sub01 Nishimoto[28] 0.658±$\pm 0.04$ —— —— 0.307±$\pm 0.14$ 11.711±$\pm 2.12$ 0.649±$\pm 0.12$ ——
Wang[32] —— —— —— 0.154±$\pm 0.11$ 13.200±$\pm 4.01$ —— ——
Ours 0.819±$\pm 0.03$ 0.783±$\pm 0.03$ 0.613±$\pm 0.07$ 0.325±$\pm 0.10$ 10.757±$\pm 2.40$ 0.820±$\pm 0.13$ 0.476±$\pm 0.41$
Sub02 Nishimoto[28] 0.661±$\pm 0.04$ —— —— 0.338±$\pm 0.14$ 11.249±$\pm 2.17$ 0.643±$\pm 0.12$ ——
Wang[32] —— —— —— 0.178±$\pm 0.11$ 13.700±$\pm 5.00$ —— ——
Ours 0.756±$\pm 0.03$ 0.759±$\pm 0.04$ 0.609±$\pm 0.07$ 0.371±$\pm 0.12$ 11.894±$\pm 2.66$ 0.834±$\pm 0.14$ 0.485±$\pm 0.44$
Sub03 Nishimoto[28] 0.657±$\pm 0.04$ —— —— 0.319±$\pm 0.14$ 10.988±$\pm 2.14$ 0.645±$\pm 0.13$ ——
Wang[32] —— —— —— 0.147±$\pm 0.11$ 12.200±$\pm 4.00$ —— ——
Ours 0.781±$\pm 0.03$ 0.793±$\pm 0.03$ 0.634±$\pm 0.07$ 0.336±$\pm 0.12$ 11.018±$\pm 2.62$ 0.834±$\pm 0.14$ 0.573±$\pm 0.39$
Table 7: Quantitative comparison of reconstruction results on HCP dataset.
E.3 Further Results on the Algonauts2021 Dataset
Figure 13: The reconstruction results on four subjects from the Algonauts2021 dataset.
Sub ID Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Sub01 Nishimoto[28] 0.688±$\pm 0.04$ —— —— 0.446±$\pm 0.08$ 9.626±$\pm 1.61$ 0.672±$\pm 0.15$ ——
Ours 0.722±$\pm 0.04$ 0.790±$\pm 0.03$ 0.599±$\pm 0.07$ 0.401±$\pm 0.10$ 10.088±$\pm 1.56$ 0.824±$\pm 0.09$ 0.439±$\pm 0.42$
Sub02 Nishimoto[28] 0.682±$\pm 0.04$ —— —— 0.443±$\pm 0.09$ 9.553±$\pm 1.53$ 0.676±$\pm 0.13$ ——
Ours 0.734±$\pm 0.04$ 0.765±$\pm 0.03$ 0.596±$\pm 0.07$ 0.465±$\pm 0.11$ 10.932±$\pm 1.68$ 0.796±$\pm 0.12$ 0.425±$\pm 0.43$
Sub03 Nishimoto[28] 0.701±$\pm 0.04$ —— —— 0.441±$\pm 0.08$ 9.576±$\pm 1.74$ 0.682±$\pm 0.16$ ——
Ours 0.679±$\pm 0.04$ 0.794±$\pm 0.03$ 0.591±$\pm 0.07$ 0.466±$\pm 0.12$ 11.089±$\pm 1.65$ 0.863±$\pm 0.11$ 0.397±$\pm 0.43$
Sub04 Nishimoto[28] 0.702±$\pm 0.04$ —— —— 0.446±$\pm 0.09$ 9.537±$\pm 1.61$ 0.665±$\pm 0.16$ ——
Ours 0.673±$\pm 0.04$ 0.810±$\pm 0.03$ 0.587±$\pm 0.07$ 0.479±$\pm 0.13$ 11.410±$\pm 1.81$ 0.848±$\pm 0.11$ 0.381±$\pm 0.42$
Sub05 Nishimoto[28] 0.676±$\pm 0.04$ —— —— 0.442±$\pm 0.08$ 9.498±$\pm 1.39$ 0.650±$\pm 0.14$ ——
Ours 0.689±$\pm 0.04$ 0.810±$\pm 0.03$ 0.592±$\pm 0.07$ 0.458±$\pm 0.12$ 10.814±$\pm 1.60$ 0.807±$\pm 0.12$ 0.406±$\pm 0.44$
Sub06 Nishimoto[28] 0.694±$\pm 0.04$ —— —— 0.444±$\pm 0.08$ 9.526±$\pm 1.57$ 0.665±$\pm 0.14$ ——
Ours 0.709±$\pm 0.04$ 0.783±$\pm 0.03$ 0.597±$\pm 0.06$ 0.489±$\pm 0.13$ 11.337±$\pm 1.70$ 0.834±$\pm 0.12$ 0.446±$\pm 0.43$
Sub07 Nishimoto[28] 0.674±$\pm 0.04$ —— —— 0.446±$\pm 0.08$ 9.630±$\pm 1.57$ 0.672±$\pm 0.14$ ——
Ours 0.681±$\pm 0.04$ 0.802±$\pm 0.03$ 0.578±$\pm 0.07$ 0.458±$\pm 0.11$ 10.889±$\pm 1.58$ 0.857±$\pm 0.11$ 0.329±$\pm 0.42$
Sub08 Nishimoto[28] 0.696±$\pm 0.04$ —— —— 0.444±$\pm 0.08$ 9.664±$\pm 1.57$ 0.662±$\pm 0.15$ ——
Ours 0.709±$\pm 0.04$ 0.802±$\pm 0.03$ 0.592±$\pm 0.07$ 0.467±$\pm 0.13$ 10.893±$\pm 1.56$ 0.820±$\pm 0.13$ 0.376±$\pm 0.43$
Sub09 Nishimoto[28] 0.673±$\pm 0.04$ —— —— 0.445±$\pm 0.08$ 9.573±$\pm 1.67$ 0.661±$\pm 0.14$ ——
Ours 0.731±$\pm 0.04$ 0.788±$\pm 0.03$ 0.594±$\pm 0.06$ 0.502±$\pm 0.12$ 11.310±$\pm 1.57$ 0.820±$\pm 0.12$ 0.400±$\pm 0.42$
Sub10 Nishimoto[28] 0.685±$\pm 0.04$ —— —— 0.441±$\pm 0.08$ 9.598±$\pm 1.79$ 0.662±$\pm 0.14$ ——
Ours 0.684±$\pm 0.04$ 0.777±$\pm 0.03$ 0.590±$\pm 0.07$ 0.465±$\pm 0.13$ 11.128±$\pm 1.91$ 0.858±$\pm 0.12$ 0.408±$\pm 0.43$
Table 8: Quantitative comparison of reconstruction results on Algonauts2021 dataset.
We conduct experiments on the Algonauts2021 competition dataset with 10 subjects, employing the same parameter settings for Mind-Animator during both training and testing as those used on the CC2017 dataset. Since no video reconstruction results has previously been published on this dataset, we replicate Nishimoto et al.’s model as a baseline and utilize video clips from the training sets of both CC2017 and HCP as video priors. Visual comparisons and quantitative assessments across various metrics are depicted in Figure 13 and detailed in Table 8, respectively. Although Table 8 indicates that our model outperforms the earlier baseline in nearly all metrics across the 10 subjects, the reconstructed results presented in Figure 13 are not entirely satisfactory, with some video frames semantically misaligned with the stimulus video. We attribute this to the scarcity of training data, which renders the video reconstruction task on this dataset challenging, given that the data volume per subject is approximately one-fifth of that in the CC2017 dataset.
Based on the experimental results across multiple datasets, we can draw the following conclusions: (1) The volume of training data from a single subject significantly influences the performance of current video reconstruction models, with greater sample size and data diversity leading to better reconstruction performance. (2) There is an urgent need to develop a new model using incremental learning or cross-subject learning methods that can be trained using data collected from different subjects, which we consider as a direction for future research.
E.4 A Detailed Ablation Study on Various Hyperparameters
[…]
Semantic-$\uparrow$
2-way-I 2-way-V VIFI-score
$\alpha$=0
(w/o fMRI-T)
0.794±$\pm 0.03$ 0.833±$\pm 0.03$ 0.594±$\pm 0.07$
$\alpha$=0.25 0.792±$\pm 0.03$ 0.823±$\pm 0.03$ 0.593±$\pm 0.08$
$\alpha$=0.5
(Our Model)
0.809±$\pm 0.03$ 0.837±$\pm 0.02$ 0.602±$\pm 0.07$
$\alpha$=0.75 0.791±$\pm 0.03$ 0.832±$\pm 0.03$ 0.594±$\pm 0.08$
$\alpha$=1.0
(w/o fMRI-V)
0.787±$\pm 0.03$ 0.812±$\pm 0.03$ 0.584±$\pm 0.08$
Table 9: Ablation study on multimodal contrastive learning. Colors reflect statistical significance (paired t-test) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
[…]
Pixel-$\uparrow$
SSIM PSNR Hue-pcc
Ratio=0 0.297±$\pm 0.11$ 9.037±$\pm 1.42$ 0.758±$\pm 0.11$
Ratio=0.2 0.276±$\pm 0.11$ 8.847±$\pm 1.59$ 0.767±$\pm 0.12$
Ratio=0.4 0.285±$\pm 0.10$ 9.045±$\pm 1.48$ 0.768±$\pm 0.11$
Ratio=0.6
(Our Model)
0.301±$\pm 0.10$ 9.134±$\pm 1.51$ 0.768±$\pm 0.11$
Ratio=0.8 0.296±$\pm 0.12$ 9.057±$\pm 1.60$ 0.767±$\pm 0.12$
Table 10: Ablation study on sparse causal mask ratio. Colors reflect statistical significance (paired t-test) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
During the training of Semantic Decoder, we control the weighting of the contrastive learning loss between fMRI-text ( $L_{BiInfoNCE}(f,t)$ ) and fMRI-video ( $L_{BiInfoNCE}(f,v)$ ) through the hyperparameter $\alpha$. We set different values for $\alpha$ (0, 0.25, 0.5, 0.75, 1.0), where $\alpha$=0 signifies the exclusion of $L_{BiInfoNCE}(f,t)$, and $\alpha$=1 signifies the exclusion of $L_{BiInfoNCE}(f,v)$.
Table 10 indicates that despite achieving optimal results for the three semantic-level metrics when $\alpha$ is set to 0.5, variations in $\alpha$ do not significantly affect the results, except when $\alpha$ is 1, suggesting that the contrastive learning loss of fMRI-video predominates. This finding aligns with the experimental results of Lin et al [16] . We attribute this phenomenon to two primary reasons: (1) The text labels are generated by the BLIP2 model, which is not a video captioning model and may suffer from "hallucination" issues, thus the labels do not perfectly reflect the video content. For instance, manual inspection reveals that the model would repetitively output "There are 1,2,3,4,5,6,7,8,9,1,2,3,4,5…" when describing a manual telephone. (2) Research by Zhang et al. [76] has shown that the effective marked length of the text branch in CLIP is relatively short, which limits its ability to process detailed descriptions.
Based on the experimental results of Tong et al.[52]  , there is significant visual information redundancy in videos that can be effectively harnessed by applying a considerable mask ratio (e.g., 0.9) . We design a sparse causal mask within the Consistency Motion Generator, which randomly masks a portion of frame tokens at a certain ratio, ensuring that the model does not take shortcuts during training. It is important to note that since we have downsampled the video to 4Hz, applying a very large mask ratio may prevent the model from learning spatiotemporal information from coherent frames. We set multiple values for the mask ratio (0, 0.2, 0.4, 0.6, 0.8) and calculate the results on pixel-level metrics, as shown in Table 10. The results indicate that the choice of mask ratio does not significantly affect the hue of reconstructed videos, but when the ratio is set to 0.6, it is significantly optimal for the other two metrics.
E.5 Further Results on Interpretability Analysis
Figure 14: Voxel-wise importance maps projected onto the visual cortex of subject 2. The lighter the color, the greater the weight of the voxel in the interpretation of feature.
Figure 15: ROI-wise importance maps in the visual cortex of subject 2.
Figure 16: Voxel-wise importance maps projected onto the visual cortex of subject 3. The lighter the color, the greater the weight of the voxel in the interpretation of feature.
Figure 17: ROI-wise importance maps in the visual cortex of subject 2.
The visualization of voxel-wise and ROI-wise importance maps on Subject 2 and 3 is depicted in the aforementioned Figures.
Appendix F Limitations and Future Work
Our model, trained on data from a single subject, is not readily generalizable to other subjects and is markedly limited by the quantity of the training data. Going forward, we intend to create a multi-subject model for brain signal decoding and video reconstruction that effectively harnesses information from a diverse array of subjects and datasets.
Appendix G Broader Impacts
The pursuit to unravel and emulate the brain’s intricate visual processing systems has been a cornerstone endeavor for scholars in computational neuroscience and artificial intelligence, aiming to catalyze the evolution of brain-inspired computational models. Recent advancements in neural decoding and the reconstruction of visual stimuli from brain activity have unveiled a lot of possibilities, yet these are accompanied by profound implications that necessitate deliberate and nuanced contemplation.
A paramount concern in this domain is the privacy and confidentiality of the neural data gathered. The integrity and security of such sensitive information are of the essence, as any compromise could have severe repercussions for individuals’ personal and medical confidentiality. This underscores an indispensable requirement for robust data protection protocols to ensure the sanctity of the data and the trust of its subjects. The onus is on the research community to implement stringent safeguards, not only to avert unauthorized access but also to foster an ethical framework that respects and protects the privacy rights of all stakeholders involved.